OpenStudy (anonymous):

A beetle sits at the top of a bicycle wheel and flies away just before it would be squashed. Assuming that the wheel turns clockwise, the beetle’s angular displacement is  rad, which corresponds to an arc length of 1.2 m. What is the wheel’s radius? Answer in units of m Leave numbers as is and please don't change due to sig figs

OpenStudy (shiraz14):

Use the formula S=r*(theta), where S is the arc length in metres (i.e. 1.2m in this context), r is the wheel radius in metres and theta is the angular displacement (in radians).

OpenStudy (anonymous):

so 1.2=r*(theta). how do I use theta?

OpenStudy (shiraz14):

Do you have the angular displacement provided in the question? I presumed you might have left it out when typing the question ...

OpenStudy (anonymous):

it is pi

OpenStudy (shiraz14):

Yes, that's right. Anyway, from the context provided in the question, we can assume that the beetle gets squashed when it lands under the wheel, implying an angular displacement of pi radians. So, the equation becomes 1.2 = pi*r, or r=(\[\frac{ 1.2 }{ \pi }\]) m (no change in sig figs).

OpenStudy (anonymous):

is it 2.618?

OpenStudy (shiraz14):

It's 0.382m.

OpenStudy (anonymous):

Thanks a lot!

OpenStudy (shiraz14):

You're most welcome.